# Brain-inspired and Self-based Artificial Intelligence

### v1

@@ 3.4 Social Self
Theory of Mind/Cognitive Empathy. Theory of Mind or Cognitive Empathy refers to a higher cognitive ability of individuals to understand their own and others’ mental states, including intentions, expectations, thoughts, and beliefs, and use this information to predict and explain the behavior of others  Premack and Woodruff (1978). One of the key milestones in the development of ToM is the acquisition of the ability to attribute false beliefs, recognizing that others may have different beliefs about the world  Wimmer and Perner (1983). This ability requires the agent to first distinguish itself from others Zeng et al. (2018b), and based on this, predict the results of others’ perception, and infer their beliefs and behaviors based on their own experience. This process closely relies on the Self’s Level 0 and Level 1 to understand one’s own perception, and the Level 2 Autonomous Self to accumulate self-experiences to distinguish self from others.
Drawing on research findings from cognitive psychology, neuroscience, and neuroimaging in the field of ToM, we integrate the multi-scale neural foundations of the theory of mind, including brain regions, brain functions, and neural circuits. We propose and construct a brain-inspired theory of mind SNN model Zeng et al. (2020). The model allows robots to learn from their own experiences and use them to infer the beliefs of others, as well as predict their behavior based on their beliefs. This enables robots to acquire the ability to think critically through false belief tasks, as shown in Figure 7a.
@@
In addition to conducting robotics experiments, we designed a virtual environment with perspective occlusion inspired by false belief tasks Baron-Cohen et al. (1985). From Figure 7b, agents with obstructed vision are unable to avoid unknown safety risks while moving. However, other agents with ToM model are able to infer dangers they cannot see through perspective transformation, thus helping them to avoid danger. Besides, in multi-agent cooperative tasks, agents with ToM model are able to infer others’ strategies based on self-experience or directly others-modelling. The inferred behavior of others influences one’s own decisions to improve the efficiency of collaboration between agents (Figure 7c).
Affective Empathy. Affective empathy is the ability to perceive and understand the feelings or emotional states of others and is the motivation for altruistic behaviors such as comforting and rescuing Waal and Preston (2017). It reflects the Social Self to a certain extent. The mirror neuron system Rizzolatti and Sinigaglia (2016), as the core of affective empathy, empathizes the others’ emotions by linking others emotional overt actions with their own emotions. The prerequisite for empathy is the self-emotional experience acquired at Level 2 Autonomous Self, which is the correlation between one’s own experienced emotions and overt actions. Therefore, seeing similar others’ overt actions triggers one’s own previous emotional response.
Our proposed BriSe AI integrated a brain-inspired affective empathy computational model based on the mirror neuron system, empowering agents to empathize with others’ emotions based on their own emotional experiences and to distinguish between self and others Feng et al. (2022). As shown in Figure 7d, we designed an affective empathy task where agents first experience the generation of their own emotions and emotional overt actions in an environment. Subsequently, they use their own experiences to empathize with their peers’ emotions and carry out rescue behaviors. The experimental results show that when perceiving the peer’s emotional overt action, it activates the same mirror neurons and emotional neurons as when the agent experiences its own emotion, realizing the sharing of emotion. Empathy for others’ negative emotions becomes an intrinsic drive for altruism, motivating the agent to sacrifice one’s own task to help others avoid negative emotions.
@@ 4.1 Fundamental Learning Strategies
Unsupervised Learning. BriSe AI-integrated unsupervised learning is based on Spike-Timing-Dependent Plasticity (STDP), and introduces adaptive synaptic filters for regulating regulate input currents, adaptive lateral inhibitory connections to ensure diverse feature extraction, and adaptive threshold balancing for maintaining the input-output balance across neural layers Dong et al. (2023). Together, these adaptation mechanisms improve the network’s information processing, learning efficiency, and feature representation capabilities.
Supervised Learning. Traditional SNN training algorithms mainly focus on the guiding role of label information in predictions, often neglecting the variability in the network’s output distribution across different time steps. This variability can lead to conflicts in the direction of optimization during the training process, thereby hindering effective optimization of the network. Similarly, during the testing process, differences in distribution between time steps can affect the network’s performance due to inaccurate output results. Our supervised learning algorithm Zhao et al. (2023a) enhances the consistency across different time steps, thereby improving the stability of SNN training and significantly enhancing performance under low-latency conditions. This algorithm combines traditional cross-entropy loss with a novel enhanced temporal consistency loss, the latter reducing differences between time steps through Kullback-Leibler divergence.
Reinforcement Learning. We adopt a layered membrane potential regularization method, and construct an efficient brain-inspired spiking neural network model for deep reinforcement learning with efficient spike information transmission Sun et al. (2022). Inspired by the brain’s state value distribution decision-making mechanism, we employ neuronal population coding to enhance the representational capability of spike sequence information, creating Gaussian neuron populations for the dimensional expansion and separable representation of binary spike sequences corresponding to continuous quantile values. Drawing inspiration from the structure of brain pyramidal neurons, we use multi-compartment neural networks to address the fusion of decision-making information from diverse sources  Sun et al. (2023). The population neuron coding approach and multi-compartment neuron information integration structure are more effective at handling multi-source decision information than the commonly used point neuron models and membrane potential computation methods in spiking neural networks.
@@
Few-shot Learning. We develop a neuromorphic dataset that reflects the complexities of human visual perception and cognition, N-Omniglot Li et al. (2022). The dataset comprises an extensive array of handwritten characters, each category represented by a minimal number of examples. In addition, we have developed a variety of few-shot learning methods, including improved nearest neighbor, convolutional network, SiameseNet, and meta-learning algorithm, demonstrate the versatility and efficacy of N-Omniglot. These methods are carefully tailored to exploit the unique characteristics of the N-Omniglot dataset, exemplify BriSe AI’s multifaceted approach to tackling complex AI challenges.
@@
Multi-modal Concept Learning. We develop a computational model for concept learning that accurately emulates the architectural paradigm of human cognitionWang and Zeng (2022a)Wang and Zeng (2024). The tripartite structure of the computational model - consisting of a multisensory information processing module, a text-derived information processing module, and a self-based semantic control module - mirrors the three networks of human cognition: a multimodal experiential system, a language-based semantic system, and a semantic control system. The main idea of our human-like concept learning is that despite the myriad of external inputs encountered by the human brain during the learning of new concepts, these inputs are all processed as unified signal transmissions within the brain. We convert the two types of concept representations into discrete neural stimuli, transforming the original representation into a spike train, thereby standardizing the data representation and simulating the initial process of concept acquisition in the human brain. What’s more, we innovatively leverage the properties of spiking neural networksWang and Zeng (2022b). The information in the temporal domain is harmonized by controlling the duration T of the performance of neurons simulated by external information, and the information from the two representations is integrated via sliding coordination. The final human-like concept representation is derived once the concept representation has been coordinated both spatially and temporally.
@@
Continual Learning. The human brain possesses the ability to learn multiple cognitive tasks sequentially and its structural developmental mechanisms allow the nervous system to dynamically expand and contract to flexibly invoke task-specific neural circuits. Motivated by this, a Dynamic Structure Development of Spiking Neural Networks (DSD-SNN) is proposed for efficient and adaptive continual learning Han et al. (2023). When a new task is coming, the DSD-SNN model first randomly grows a part of neurons forming a new pathway to learn new knowledge and receive old knowledge that can be reused. During training for new tasks, the DSD-SNN gradually prune the redundant neurons that are continuously inactive for several epochs. Dynamic neuronal expansion and pruning increase memory capacity and reduce energy consumption, and the overlapping shared structure helps to quickly leverage all acquired knowledge to new tasks.
@@
where $\eta_{t}$ denotes the learnable coefficient at time step $t$ and $\sigma$ represents the sigmoid function. $\mathop{CKA^{\prime}}$ represents the computation of the kernel function of the vectors followed by the computation of central kernel alignment (CKA) Kornblith et al. (2019). $g\left(\mathbf{x}_{s}^{i},t\right)$ and $g\left(\mathbf{x}_{t}^{j},t\right)$ denote the features of the data in the source and target domains respectively after the encoder $g$. $\ell_{cls-e}$ is classification loss of event data.
Evolutionary Learning. For evolutionary learning, the NeuEvo Shen et al. (2023) framework introduces an integrated approach for evolving the neural circuitry of SNNs, inspired by the dynamic and adaptive nature of biological neural systems. This integrated approach is not solely focused on evolutionary computation, but rather encompasses a comprehensive strategy that blends local synaptic dynamics with global network optimization. The framework leverages a synergistic process for synaptic weight adaptation. This process combines local synaptic activity, governed by the STDP rule, with global error feedback for network-wide optimization. The synaptic weight ($w$) at any given neuron is updated based on the following formulation:
@@
In addition, the ELSM Pan et al. (2024) and MSE-NAS Pan et al. (2023a) specially design adaptive structure evolution for recurrent and deep spiking neural networks, respectively. Through exploring the small-world characteristics that exist in the human brain, ELSM guides the evolution of network topology from both static and dynamic perspectives. MSE-NAS evolves microscopic neuron computations, mesoscopic microcircuit combinations, and macroscopic global brain area connectivity, supplemented by customised brain-inspired evaluation function, encoding scheme, and genetic operations. Adaptive LSM Pan et al. (2023b) deeply draws inspiration from brain evolution mechanisms, integrating structural evolution with multi-scale biological plasticity learning rules.
@@ 4.2 Self-organized Learning supported Self Implementation
As a neurotransmitter, dopamine plays a critical role in emotions, memory, action selection, motivation, attention, reward-driven behavior, and motor control Wise (2004), and thus provides constructive inspiration for the development of the autonomous self. The neuromodulatory role of the dopamine system is also responsible for trial-and-error learning in decision-making, whereby the reward prediction error signal carried by dopamine directly modulates the connection weights from the prefrontal cortex to the striatum. Specifically, global long-term neuromodulation enables the responsiveness of local synaptic plasticity, yielding dopamine-regulated local synaptic plasticity. In this paper, We adopt the reward-modulated STDP Izhikevich (2007) that is capable of solving the problem of credit assignment.
@@
Utilizing the inhibitory control mechanism of a single neuron as exemplified in our false belief task Zeng et al. (2020), the Inferior Parietal Lobule (IPL) or ventromedial Prefrontal Cortex (vmPFC) receive currents related to both self-relevant and other-relevant stimuli. The currents feeding into inhibitory and temporary neurons vary depending on whether the inference involves others’ beliefs (encompassing self-perspective and self-belief inhibition) or self-beliefs (including other-perspective and other-belief inhibition). During self-perspective and self-belief inhibition phases, the currents to inhibitory neurons align with other-related stimuli, whereas currents to temporary neurons correspond to self-related stimuli. Consequently, other-related stimuli are initially processed, followed by self-related stimuli. Conversely, in the other-perspective and other-belief inhibition phases, inhibitory neurons exhibit intense activity, capable of completely suppressing other-relevant stimuli, while inputs to temporary neurons mirror other-related stimuli. Thus, self-related stimuli precede processing, followed by other-related stimuli.
@@
Self-experience inferring others. In the context of ToM and affective empathy, self-experience aids in inferring others’ mental states and empathizing with others’ emotions. This concept is also referred to as ’self-projection’ Buckner and Carroll (2007), involves using personal memories to understand others’ perspectives Moreau et al. (2013). Self-experience is derived from the interaction between the decision model and the environment. When constructing the ToM model, we incorporated a buffer for storing self-experience and a sub-module for inferring others Zhao et al. (2023c). A SNN-based sub-module was developed to simulate the mapping from self-experience learning to inferring others’ behaviors. By using observations of others as input, our model can predict their behaviors.
Mirror neuron system for affective empathy. Psychological research suggests that observers can use the perception-action mechanism (PAM) to empathize with others Preston and de Waal (2017). Perceiving another person’s emotional state activates the same emotional representation in the observer’s brain, which is equivalent to the observer also experiencing the emotion, and the mirror neuron system (MNS) provides the biological basis for PAM Erhan et al. (2006); Carr et al. (2003). The mirror neuron system contains mirror neurons and anti-mirror neurons. Mirror neurons enable individuals to empathize with others’ emotions through their own emotional experiences, while anti-mirror neurons help distinguish the originator of the emotion, reflecting primary self-awareness Christian and Valeria (2010). Our proposed mirror neuron system-based affective empathy SNN contains three clusters of neuronal populations: the emotional cortex representing different emotional states, the motor cortex containing the mirror neuron system, and the perceptual cortex representing emotional overt actions. These modules are interconnected with the mirror neuron system as the core, and local synaptic plasticity is used to learn the connection strength, achieving empathy for others’ emotions. Additionally, we replicate the emergence of anti-mirror neurons, achieving self-other differentiation.

### v2

@@ 3.4 Social Self
Theory of Mind/Cognitive Empathy. Theory of Mind or Cognitive Empathy refers to a higher cognitive ability of individuals to understand their own and others’ mental states, including intentions, expectations, thoughts, and beliefs, and use this information to predict and explain the behavior of others  [38]. One of the key milestones in the development of ToM is the acquisition of the ability to attribute false beliefs, recognizing that others may have different beliefs about the world  [39]. This ability requires the agent to first distinguish itself from others [40], and based on this, predict the results of others’ perception, and infer their beliefs and behaviors based on their own experience. This process closely relies on the Self’s Level 0 and Level 1 to understand one’s own perception, and the Level 2 Autonomous Self to accumulate self-experiences to distinguish self from others.
Drawing on research findings from cognitive psychology, neuroscience, and neuroimaging in the field of ToM, we integrate the multi-scale neural foundations of the theory of mind, including brain regions, brain functions, and neural circuits. We propose and construct a brain-inspired theory of mind SNN model [41]. The model allows robots to learn from their own experiences and use them to infer the beliefs of others, as well as predict their behavior based on their beliefs. This enables robots to acquire the ability to think critically through false belief tasks, as shown in Figure 7a.
@@
In addition to conducting robotics experiments, we designed a virtual environment with perspective occlusion inspired by false belief tasks [42]. From Figure 7b, agents with obstructed vision are unable to avoid unknown safety risks while moving. However, other agents with ToM model are able to infer dangers they cannot see through perspective transformation, thus helping them to avoid danger. Besides, in multi-agent cooperative tasks, agents with ToM model are able to infer others’ strategies based on self-experience or directly others-modelling. The inferred behavior of others influences one’s own decisions to improve the efficiency of collaboration between agents (Figure 7c).
Affective Empathy. Affective empathy is the ability to perceive and understand the feelings or emotional states of others and is the motivation for altruistic behaviors such as comforting and rescuing [43]. It reflects the Social Self to a certain extent. The mirror neuron system [44], as the core of affective empathy, empathizes the others’ emotions by linking others emotional overt actions with their own emotions. The prerequisite for empathy is the self-emotional experience acquired at Level 2 Autonomous Self, which is the correlation between one’s own experienced emotions and overt actions. Therefore, seeing similar others’ overt actions triggers one’s own previous emotional response.
Our proposed BriSe AI integrated a brain-inspired affective empathy computational model based on the mirror neuron system, empowering agents to empathize with others’ emotions based on their own emotional experiences and to distinguish between self and others [45]. As shown in Figure 7d, we designed an affective empathy task where agents first experience the generation of their own emotions and emotional overt actions in an environment. Subsequently, they use their own experiences to empathize with their peers’ emotions and carry out rescue behaviors. The experimental results show that when perceiving the peer’s emotional overt action, it activates the same mirror neurons and emotional neurons as when the agent experiences its own emotion, realizing the sharing of emotion. Empathy for others’ negative emotions becomes an intrinsic drive for altruism, motivating the agent to sacrifice one’s own task to help others avoid negative emotions.
@@ 4.1 Fundamental Learning Strategies
Unsupervised Learning. BriSe AI-integrated unsupervised learning is based on Spike-Timing-Dependent Plasticity (STDP), and introduces adaptive synaptic filters for regulating regulate input currents, adaptive lateral inhibitory connections to ensure diverse feature extraction, and adaptive threshold balancing for maintaining the input-output balance across neural layers [19]. Together, these adaptation mechanisms improve the network’s information processing, learning efficiency, and feature representation capabilities.
Supervised Learning. Traditional SNN training algorithms mainly focus on the guiding role of label information in predictions, often neglecting the variability in the network’s output distribution across different time steps. This variability can lead to conflicts in the direction of optimization during the training process, thereby hindering effective optimization of the network. Similarly, during the testing process, differences in distribution between time steps can affect the network’s performance due to inaccurate output results. Our supervised learning algorithm [20] enhances the consistency across different time steps, thereby improving the stability of SNN training and significantly enhancing performance under low-latency conditions. This algorithm combines traditional cross-entropy loss with a novel enhanced temporal consistency loss, the latter reducing differences between time steps through Kullback-Leibler divergence.
Reinforcement Learning. We adopt a layered membrane potential regularization method, and construct an efficient brain-inspired spiking neural network model for deep reinforcement learning with efficient spike information transmission [35]. Inspired by the brain’s state value distribution decision-making mechanism, we employ neuronal population coding to enhance the representational capability of spike sequence information, creating Gaussian neuron populations for the dimensional expansion and separable representation of binary spike sequences corresponding to continuous quantile values. Drawing inspiration from the structure of brain pyramidal neurons, we use multi-compartment neural networks to address the fusion of decision-making information from diverse sources  [46]. The population neuron coding approach and multi-compartment neuron information integration structure are more effective at handling multi-source decision information than the commonly used point neuron models and membrane potential computation methods in spiking neural networks.
@@
Few-shot Learning. We develop a neuromorphic dataset that reflects the complexities of human visual perception and cognition, N-Omniglot [21]. The dataset comprises an extensive array of handwritten characters, each category represented by a minimal number of examples. In addition, we have developed a variety of few-shot learning methods, including improved nearest neighbor, convolutional network, SiameseNet, and meta-learning algorithm, demonstrate the versatility and efficacy of N-Omniglot. These methods are carefully tailored to exploit the unique characteristics of the N-Omniglot dataset, exemplify BriSe AI’s multifaceted approach to tackling complex AI challenges.
@@
Multi-modal Concept Learning. We develop a computational model for concept learning that accurately emulates the architectural paradigm of human cognition[47][48]. The tripartite structure of the computational model - consisting of a multisensory information processing module, a text-derived information processing module, and a self-based semantic control module - mirrors the three networks of human cognition: a multimodal experiential system, a language-based semantic system, and a semantic control system. The main idea of our human-like concept learning is that despite the myriad of external inputs encountered by the human brain during the learning of new concepts, these inputs are all processed as unified signal transmissions within the brain. We convert the two types of concept representations into discrete neural stimuli, transforming the original representation into a spike train, thereby standardizing the data representation and simulating the initial process of concept acquisition in the human brain. What’s more, we innovatively leverage the properties of spiking neural networks[49]. The information in the temporal domain is harmonized by controlling the duration T of the performance of neurons simulated by external information, and the information from the two representations is integrated via sliding coordination. The final human-like concept representation is derived once the concept representation has been coordinated both spatially and temporally.
@@
Continual Learning. The human brain possesses the ability to learn multiple cognitive tasks sequentially and its structural developmental mechanisms allow the nervous system to dynamically expand and contract to flexibly invoke task-specific neural circuits. Motivated by this, a Dynamic Structure Development of Spiking Neural Networks (DSD-SNN) is proposed for efficient and adaptive continual learning [22]. When a new task is coming, the DSD-SNN model first randomly grows a part of neurons forming a new pathway to learn new knowledge and receive old knowledge that can be reused. During training for new tasks, the DSD-SNN gradually prune the redundant neurons that are continuously inactive for several epochs. Dynamic neuronal expansion and pruning increase memory capacity and reduce energy consumption, and the overlapping shared structure helps to quickly leverage all acquired knowledge to new tasks.
@@
where $\eta_{t}$ denotes the learnable coefficient at time step $t$ and $\sigma$ represents the sigmoid function. $\mathop{CKA^{\prime}}$ represents the computation of the kernel function of the vectors followed by the computation of central kernel alignment (CKA) [50]. $g\left(\mathbf{x}_{s}^{i},t\right)$ and $g\left(\mathbf{x}_{t}^{j},t\right)$ denote the features of the data in the source and target domains respectively after the encoder $g$. $\ell_{cls-e}$ is classification loss of event data.
Evolutionary Learning. For evolutionary learning, the NeuEvo [24] framework introduces an integrated approach for evolving the neural circuitry of SNNs, inspired by the dynamic and adaptive nature of biological neural systems. This integrated approach is not solely focused on evolutionary computation, but rather encompasses a comprehensive strategy that blends local synaptic dynamics with global network optimization. The framework leverages a synergistic process for synaptic weight adaptation. This process combines local synaptic activity, governed by the STDP rule, with global error feedback for network-wide optimization. The synaptic weight ($w$) at any given neuron is updated based on the following formulation:
@@
In addition, the ELSM [51] and MSE-NAS [52] specially design adaptive structure evolution for recurrent and deep spiking neural networks, respectively. Through exploring the small-world characteristics that exist in the human brain, ELSM guides the evolution of network topology from both static and dynamic perspectives. MSE-NAS evolves microscopic neuron computations, mesoscopic microcircuit combinations, and macroscopic global brain area connectivity, supplemented by customised brain-inspired evaluation function, encoding scheme, and genetic operations. Adaptive LSM [53] deeply draws inspiration from brain evolution mechanisms, integrating structural evolution with multi-scale biological plasticity learning rules.
@@ 4.2 Self-organized Learning supported Self Implementation
As a neurotransmitter, dopamine plays a critical role in emotions, memory, action selection, motivation, attention, reward-driven behavior, and motor control [54], and thus provides constructive inspiration for the development of the autonomous self. The neuromodulatory role of the dopamine system is also responsible for trial-and-error learning in decision-making, whereby the reward prediction error signal carried by dopamine directly modulates the connection weights from the prefrontal cortex to the striatum. Specifically, global long-term neuromodulation enables the responsiveness of local synaptic plasticity, yielding dopamine-regulated local synaptic plasticity. In this paper, We adopt the reward-modulated STDP [55] that is capable of solving the problem of credit assignment.
@@
Utilizing the inhibitory control mechanism of a single neuron as exemplified in our false belief task [41], the Inferior Parietal Lobule (IPL) or ventromedial Prefrontal Cortex (vmPFC) receive currents related to both self-relevant and other-relevant stimuli. The currents feeding into inhibitory and temporary neurons vary depending on whether the inference involves others’ beliefs (encompassing self-perspective and self-belief inhibition) or self-beliefs (including other-perspective and other-belief inhibition). During self-perspective and self-belief inhibition phases, the currents to inhibitory neurons align with other-related stimuli, whereas currents to temporary neurons correspond to self-related stimuli. Consequently, other-related stimuli are initially processed, followed by self-related stimuli. Conversely, in the other-perspective and other-belief inhibition phases, inhibitory neurons exhibit intense activity, capable of completely suppressing other-relevant stimuli, while inputs to temporary neurons mirror other-related stimuli. Thus, self-related stimuli precede processing, followed by other-related stimuli.
@@
Self-experience inferring others. In the context of ToM and affective empathy, self-experience aids in inferring others’ mental states and empathizing with others’ emotions. This concept is also referred to as ’self-projection’ [56], involves using personal memories to understand others’ perspectives [57]. Self-experience is derived from the interaction between the decision model and the environment. When constructing the ToM model, we incorporated a buffer for storing self-experience and a sub-module for inferring others [58]. A SNN-based sub-module was developed to simulate the mapping from self-experience learning to inferring others’ behaviors. By using observations of others as input, our model can predict their behaviors.
Mirror neuron system for affective empathy. Psychological research suggests that observers can use the perception-action mechanism (PAM) to empathize with others [59]. Perceiving another person’s emotional state activates the same emotional representation in the observer’s brain, which is equivalent to the observer also experiencing the emotion, and the mirror neuron system (MNS) provides the biological basis for PAM [60, 61]. The mirror neuron system contains mirror neurons and anti-mirror neurons. Mirror neurons enable individuals to empathize with others’ emotions through their own emotional experiences, while anti-mirror neurons help distinguish the originator of the emotion, reflecting primary self-awareness [62]. Our proposed mirror neuron system-based affective empathy SNN contains three clusters of neuronal populations: the emotional cortex representing different emotional states, the motor cortex containing the mirror neuron system, and the perceptual cortex representing emotional overt actions. These modules are interconnected with the mirror neuron system as the core, and local synaptic plasticity is used to learn the connection strength, achieving empathy for others’ emotions. Additionally, we replicate the emergence of anti-mirror neurons, achieving self-other differentiation.